In a study to estimate the proportion of smokers among the residents in a certain city and its suburbs, it is found that 24 of 104 urban residents are smokers, while 34 of 108 suburban residents are smokers.


1
Expert's answer
2022-05-17T11:11:23-0400

The value of the pooled proportion is computed as


"\bar{p}=\dfrac{X_1+X_2}{n_1+n_2}=\dfrac{24+34}{104+108}=\dfrac{29}{106}"

The following null and alternative hypotheses for the population proportion needs to be tested:

"H_0:p_1=p_2"

"H_a:p_1\not=p_2"

This corresponds to a two-tailed test, and a z-test for two population proportions will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a two-tailed test is"z_c = 1.96."

The rejection region for this two-tailed test is "R = \{z: |z| > 1.96\}."

The z-statistic is computed as follows:


"z=\dfrac{\hat{p}_1-\hat{p}_2}{\sqrt{\bar{p}(1-\bar{p})(1/n_1+1/n_2)}}"

"=\dfrac{\dfrac{24}{104}-\dfrac{34}{108}}{\sqrt{\dfrac{29}{106}(1-\dfrac{29}{106})(1/104+1/108)}}"

"=-1.380342"

Since it is observed that "|z| = 1.380342 \le 1.96=z_c ," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value is "p=2P(Z<-1.380342)=0.167481," and since "p=0.167481>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population proportion "p_1" is different than "p_2," at the "\alpha = 0.05" significance level.